OpenStudy (anonymous):

$8,000 is invested in two different accounts earning 3% and 4% interest. If the two accounts earn a total of $295 in interest, how much money was invested at 3%? $2,500 $3,500 $5,500

OpenStudy (anonymous):

Let x be the amount invested at 3%, then \[0.04 (8000-x)+0.03 x=295 \] Solve for x

OpenStudy (anonymous):

Simplifying \[320.\, -0.01 x=295 \\ .01 x = 320 -295=25\\ x= 2500 \]

OpenStudy (inkyvoyd):

@eliassaab , why don't you let Corey work it out on his own?